OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = 6x^2 - 24x + 32?

OpenStudy (anonymous):

\[-\frac{b}{2a}\] give the first coordinate of the vertex

OpenStudy (anonymous):

\[-\frac{-24}{2\times 6}=2\]

OpenStudy (anonymous):

second coordinate is what you get when you replace x by 2

OpenStudy (anonymous):

-8?

OpenStudy (anonymous):

+8

OpenStudy (anonymous):

(2,8)

OpenStudy (anonymous):

thanks